OpenStudy (anonymous):

Find the limit. Use L'Hopital's rule if it applies. lim [1-cosh (8x)]/x^2 =??? x → 0

OpenStudy (anonymous):

any help would be appreciated!! thank you!!!

myininaya (myininaya):

so does l'hosptal apply do we have 0/0?

myininaya (myininaya):

by the way \[\cosh (8x)=\frac{e^{8x}+e^{-8x}}{2}\] if that helps you decide if l'hosptial applies

OpenStudy (anonymous):

so its 2/2?

OpenStudy (anonymous):

which is 1?

myininaya (myininaya):

and we would then have (1-1)/0^2 so we do have 0/0 :)

myininaya (myininaya):

and that means we can apply l'hospital's rule

OpenStudy (anonymous):

yes that's what i have too in one of my steps, but the limit i got once having applied the rule was 0 but my hw said to was wrong ;/

OpenStudy (anonymous):

it*

myininaya (myininaya):

do you prefer to leave cosh(8x) or can we write it in terms of those exponential functions your call i prefer the exponential function way because I'm actually less familiar with all the properties derived from the hyperbolic functions

OpenStudy (anonymous):

yeah the exponential is fine :)

myininaya (myininaya):

\[\lim_{x \rightarrow 0}\frac{1-\frac{e^{8x}+e^{-8x}}{2}}{x^2}=\lim_{x \rightarrow 0}\frac{2-(e^{8x}+e^{-8x})}{2x^2} =\lim_{x \rightarrow 0} \frac{2-e^{8x}-e^{-8x}}{2x^2}\]

myininaya (myininaya):

ok so this is the function we have

myininaya (myininaya):

the derivative of the bottom is pretty easy...

myininaya (myininaya):

so let's just jump straight to the top

OpenStudy (anonymous):

sounds great

myininaya (myininaya):

what is the derivative of the top ?

OpenStudy (anonymous):

so its -8e^8x+8e^8x?

OpenStudy (anonymous):

for the top

myininaya (myininaya):

\[\lim_{x \rightarrow 0}\frac{0-8e^{8x}+8e^{-8x}}{4x}\] well that one exponent show still be negative but other than that great :)

OpenStudy (anonymous):

oops yeah i always lose my signs ;/ lol. so now we just plug in the limit for each and see what we get?

myininaya (myininaya):

yep and to our surprise we have ?

myininaya (myininaya):

drum rolls

OpenStudy (anonymous):

the bottom limit would be 0 so wouldn't that automatically make it 0?

myininaya (myininaya):

well both top and bottom equals 0

myininaya (myininaya):

0/0 so we have to do l'hospital again

myininaya (myininaya):

but don't worry this is the last time

OpenStudy (anonymous):

lol okay lets see

myininaya (myininaya):

I know you can do the bottom

myininaya (myininaya):

try differentiating that top again

myininaya (myininaya):

well the new top we have

OpenStudy (anonymous):

okay hang on

myininaya (myininaya):

k hanging

OpenStudy (anonymous):

-64e^-8x(e^(16x)+1 :S lol

myininaya (myininaya):

ok ummm... well hmmm... let me help you with that

myininaya (myininaya):

\[(-8e^{8x}+8e^{-8x})' \\ -8(8x)'e^{8x}+8(-8x)'e^{-8x} =?\]

OpenStudy (anonymous):

-64xe^(8x)-64xe^(-8x)

myininaya (myininaya):

tons better

myininaya (myininaya):

\[\lim_{x \rightarrow 0}\frac{-64e^{8x}-64e^{-8x}}{4}\]

OpenStudy (anonymous):

lol so now i just take the limit as x-->0?

myininaya (myininaya):

see how we want use l'hosptial again because the bottom is 4

myininaya (myininaya):

right

myininaya (myininaya):

this is the final step

myininaya (myininaya):

just plug in 0

OpenStudy (anonymous):

ohhhokay that makes tons of sense

myininaya (myininaya):

won't* (not want)

OpenStudy (anonymous):

alrighty! thank you soooo much!!! you were a great help!!1

myininaya (myininaya):

thanks

myininaya (myininaya):

and np